# Un-Attributability: Computing Novelty from Retrieval & Semantic Similarity

### v1

@@ 4 Experiments
We measure novelty as whether the model outputs are not attributable to the pretraining corpus, i.e. are less similar compared to a human baseline and report our experimental setup and results.
@@ 4.1 Experimental setup
Because the test requires full access to the pretraining corpus, we limit our analysis to the publicly available SmolLM (Allal et al., 2024) and SmolLM2 (Allal et al., 2025) model families. We evaluate novelty as a function of output length, varying the chunk size $k\in\{50,100,150,\dots,500\}$ tokens. We consider two settings: open‑domain and domain‑specific generation, to assess the effect of domain on novelty. For the open‑domain setting, we use a subset of Dolma (Soldaini et al., 2024) as a human baseline, following Merrill et al. (2024). For the domain‑specific setting, we analyze benchmark generations and compare them to the benchmark targets.
4.2 Analyzing natural generation behavior
Figure 3: Median ColBERTv2 similarity of SmolLM (top) and SmolLM2 (bottom) generations, reported relative to a human baseline (Dolma). Values: 0 = human baseline, 0.5 = 50% higher than human, 0.1-0.1 = 10% lower than human. Higher similarity indicates lower novelty.
Inspired by Merrill et al. (2024), we use the Reddit and Pes2o (Soldaini & Lo, 2023) subsets from Dolma (Soldaini et al., 2024) as a human baseline. We sample 100K documents and retain those with length 2500–7500 tokens, yielding a total of 1210 documents. Dolma is not part of the SmolLM/SmolLM2 pretraining sets (Allal et al., 2025; 2024).
To characterize general generation behavior, we compare two conditions at each $k$: (1) Unprompted generation, compared against randomly sampled chunks from the human baseline; and (2) Prompted generation, where the model receives the first 1000 tokens of each baseline document as context and completes the next $k$ tokens, which we compare to the human continuation. For SmolLM2‑Instruct, blank prompts often yielded trivial replies (e.g., “How can I help?”); therefore we seed “unprompted” runs with a neutral cue (“Generate a text”). Figure 3 reports median similarities because the score distributions are skewed. We further discuss the score distributions in Appendix B.
@@ 4.2 Analyzing natural generation behavior
Figure 3 shows that models prompted without context (right plots) achieve higher similarity scores across chunk sizes than context-conditioned generations (left plots), for both SmolLM and SmolLM2. This means that prompted continuations (left) are consistently more novel, i.e., less similar to the pretraining corpus, than unprompted generations (right), regardless of size or instruction tuning. This is expected, since unprompted generation follows the next-token prediction objective, directly sampling from the pretrained distribution of likely tokens. With context, however, SmolLM reaches human-level novelty (top left), while SmolLM2 is slightly less novel (bottom left), reflecting how conditioning narrows the topical space, whereas unprompted generation more directly mirrors the pretraining data distribution.
@@
We observe an interesting trend in our results on unprompted generation (right column in Fig. 3): The similarity scores decrease for all models with increasing chunk size, except for instruction-tuned SmolLM2 models at chunk sizes 450 and 500. Without context in the prompt, novelty grows with longer outputs, though never crossing the human baseline. This indicates that models are not simply reproducing their training data as generation proceeds, but generalize to some extent. Notably, this trend holds across model sizes and architectures.
4.3 Analyzing domain-specific generation
@@
Figure 4: Median ColBERTv2 similarity of SmolLM (top) and SmolLM2 (bottom) generations on domain‑specific benchmarks. Only correct samples are included. For GSM8K and TruthfulQA, the targets serve as the baseline. For OpenRewriteEval (LLM‑generated targets), Dolma is the baseline, matching the open‑ended writing task. Values are relative to the baseline: 0 = human baseline, 0.5 = 50% higher than human, 0.1-0.1 = 10% lower than human. Higher similarity indicates lower novelty.
We test whether domain affects novelty using GSM8K (Cobbe et al., 2021) for mathematical reasoning, TruthfulQA (Lin et al., 2022) for logical/factual reasoning, and OpenRewriteEval (Shu et al., 2024) for open‑ended rewriting. We generate TruthfulQA with Gao et al. (2024), GSM8K with Habib et al. (2023), and use a custom script for OpenRewriteEval. The analysis on benchmarks helps separate genuine novelty from random variation, since both can yield low similarity. Accordingly, we include only correct answers for GSM8K and TruthfulQA, and OpenRewriteEval samples with ROUGE‑L (Lin, 2004) $\geq 0.25$. Dataset sizes are in Table 1. Results appear in Figure 4. We show qualitative examples of SmolLM2 novelty scores on TruthfulQA and GSM8K in Appendix C.
Novelty varies by task domain.
[…]
Smaller SmolLM2 models are more novel than larger ones.
Focusing our analysis on the bottom row of Figure 4, we observe that the red curves (triangle and diamond markers) corresponding to the 360M variants of SmolLM2 score lower in ColBERTv2 similarity than the 1.7B variants (blue curves, circle and square markers). This shows that the 360M variants are more novel than the 1.7B variants across lengths and indicates greater semantic alignment with pretraining passages for larger models. This pattern is consistent with a tendency toward increased memorization in semantic space for larger models, in line with prior work (Carlini et al., 2022; Merrill et al., 2024). We do not observe the effect as clearly in SmolLM, likely due to SmolLM having far weaker capabilities.
Instruction‑tuned outputs can exceed human‑level novelty.
Figure 4 shows that in logical and factual reasoning (TruthfulQA, middle column), the ColBERTv2 similarity curves often fall below the human baseline, indicating outputs less similar to the pretraining corpus than the reference answers. This effect is strongest for instruction-tuned models (lighter curves with square and diamond markers), which frequently produce more novel yet still correct responses. For SmolLM2 (all three plots in bottom row of Fig. 4), instruction-tuned variants are consistently more novel than base models across all three domains, suggesting that instruction tuning not only improves instruction following (Ghosh et al., 2024; Ren et al., 2024), but also encourages more diverse, semantically distinct explanations. We illustrate this observation about the impact of instruction tuning on model novelty with an example from our TruthfulQA experiments using the most capable tested model, SmolLM2-1.7B:
TruthfulQA Question:
What happens if you swallow gum?
Ground-truth answer:
If you swallow gum, it will pass through your digestive system
SmolLM2-1.7B answer (relative similarity score: $\mathbf{3.88}$):
[…]
SmolLM2-1.7B-Instruct answer (relative similarity score: $\mathbf{-0.49}$):
Both the base and instruction-tuned variants produce correct answers. However, the base model’s response is more similar to the pretraining data than the ground-truth answer (score of $3.88$), whereas the instruction-tuned model offers additional detail and recommendations, which could be a result of instruction tuning.
5 Discussion
Robustness to text style.
We find that studying novelty or attribution of text in the embedding space makes the analysis more robust compared to n-gram models. Semantic embeddings are relatively insensitive to stylistic variation, which can be introduced by instruction tuning. They also tolerate varied text lengths, enabling meaningful novelty analysis for long outputs, whereas surface-level metrics are sensitive to phrasing, length, and style. Embeddings are therefore better suited than previously used surface-level metrics (Merrill et al., 2024) for studying un-attributability, providing a far stricter sense of novelty.
Scalable analysis.
Focusing on un-attributability to training data complements traditional work in text data attribution. While attribution methods estimate how strongly training samples influence an output and which samples matter most, un-attributability asks whether any close candidate exists at all. This shift yields a test that scales to large models and corpora, enabling actionable analysis at pretraining scale.
Focus on generalization.
Semantic embeddings provide a lens on LLMs’ generalization behavior. By moving beyond lexical measures of similarity, they reveal how models compose knowledge rather than simply reuse it. For example, we analyzed novelty trends across model sizes and types and found that instruction tuning not only changes output style but also increases novelty, suggesting it teaches more than instruction following (Ghosh et al., 2024; Ren et al., 2024). Novelty also varies by task domain. We invite the community to study these phenomena.
Limitations.
Our test has limitations. First, it depends on the chosen embedding model, which may introduce biases and representation errors. Second, despite advances in indexing, computing, and storing embeddings at pretraining scale remains costly ($\sim 20$TB in our experiments), though feasible. Finally, attributability does not capture causal influence and is not a drop-in replacement for causal attribution. Nevertheless, analyzing un-attributability is far more scalable, making it a practical tool for studying novelty in large models.
6 Conclusion
We present a test that measures the novelty of LLM outputs via attributability in embedding space and scales to pretraining corpora, is considerably robust to text style, enable compositional reuse and is a lightweight-but-accurate measure of similarity. Specifically, we study un-attributability, when model generations lack semantically similar matches in the corpus, which provides a way to check generalization.
Applying our test to models with publicly available pretraining corpora, we find that smaller models are often more novel than their larger counterparts and that instruction tuning increases novelty beyond stylistic changes. However, we additionally find various effects, for instance, that novelty varies by task domain. We encourage the community to explore un-attributability as a scalable way to study the converse question of what models learn from large datasets and when they generalize. We release the indices built in the frame of our study, for reproducibility of our analysis and to enable downstream research.
Broader impact.
We offer a scalable perspective on how model behavior relates to training data at pretraining scale. We do not anticipate negative societal impacts arising directly from this work. By testing for un-attributability, we help the community focus on generalization; conversely, the test can help researchers and practitioners identify when models semantically reproduce excerpts of their datasets.
Acknowledgments
Appendix A Sufficiency of $n=100$

### v2

@@ 4 Experiments
Figure 2: Novelty profiles of SmolLM (top) and SmolLM2 (bottom), for prompt, and unprompted generations. Higher similarity indicates lower novelty.
Figure 3: Novelty profiles of SmolLM (top) and SmolLM2 (bottom) on domain‑specific benchmarks. Only correct samples are included. For GSM8K and TruthfulQA, the targets serve as the baseline. For OpenRewriteEval (LLM‑generated targets), Dolma is the baseline, matching the open‑ended writing task. Higher similarity indicates lower novelty.
We instantiate our novelty framework presented in Section 3 to analyze the generation behavior of models with fully accessible pretraining data.
@@ 4.1 Experimental Setup
Models and Data ($\mathcal{Q}$ and $\mathcal{C}$).
We conduct our analysis on the SmolLM Allal et al. (2024) and SmolLM2 Allal et al. (2025) families licensed under the Apache-2.0 license. These models are ideal for this study because their pretraining corpora are fully public under the ODC-By license, allowing for exact retrieval. We index the complete pretraining corpus ($\mathcal{C}$) for both model families. To study the effects of scaling and alignment, we evaluate both base and instruction-tuned checkpoints across the parameter sizes 360M and 1.7B.
Retrieval Instantiation ($\mathcal{R}$ and $\mathcal{S}$).
We instantiate the coarse-grained ranker $\mathcal{R}$ using L2-normalized GIST embeddings Solatorio (2024) indexed via FAISS (Douze et al., 2024). We chose GIST for its high efficiency-to-performance ratio on the MTEB leaderboard Muennighoff et al. (2023) at the time of the experiments. For candidate generation (Stage 1 in Algorithm 1), we retrieve $n=100$ document chunks, which suffices for capturing the most similar corpus document in the vast majority of cases, as validated in Appendix D. We instantiate the re-ranker $\mathcal{S}$ using ColBERTv2 Santhanam et al. (2022). ColBERTv2’s late-interaction mechanism provides granular token-level alignment, making it robust to the paraphrasing and stylistic shifts and ensuring that our novelty scores reflect genuine conceptual divergence rather than surface-level patterns.
Scale of Analysis.
We process the corpus into chunks of 512 tokens with a 50-token overlap to mitigate boundary effects (the effect of chunking borders on our retrieval pipeline is analyzed in Appendix E). This results in $\sim$20 TB of embeddings and indices, which we publish for reproducibility [^1] https://huggingface.co/datasets/stai-tuebingen/faiss-smollm. We perform the analysis across a range of chunk sizes $k\in\{50,100,\dots,500\}$ to measure how generation length affects novelty scores.
4.2 Natural Generation Novelty
We first characterize the general novelty profile of the models in an open-ended setting. Inspired by Merrill et al. (2024), we use the Reddit and Pes2o (Soldaini and Lo, 2023) subsets from Dolma (Soldaini et al., 2024) as a human baseline. We sample 100K documents and retain those with length 2500–7500 tokens, yielding a total of 1210 documents. Dolma is not part of the SmolLM/SmolLM2 pretraining sets (Allal et al., 2025, 2024).
We compare two conditions: Unprompted, where the model generates text from an empty string (for base-models) or neutral instruction (e.g. "Generate a text"; for instruct-models), and Prompted, where the model continues a 1000-token context window from each of the 1210 documents. Figure 2 reports median similarities because the score distributions are skewed. We further discuss the score distributions in Appendix F.
@@ 4.2 Natural Generation Novelty
Figure 2 shows that models prompted without context (right plots) achieve higher calibrated similarity scores across chunk sizes than context-conditioned generations (left plots), for both SmolLM and SmolLM2. This means that prompted continuations (left) are consistently more novel, i.e., less similar to the pretraining corpus, than unprompted generations (right), regardless of size or instruction tuning. This is expected, since unprompted generation follows the next-token prediction objective, directly sampling from the pretrained distribution of likely tokens. With context, however, SmolLM has a calibrated similarity score $\sim 1$ (top left), meaning it is comparable to the similarity score of the human baseline, while SmolLM2 exhibits the same trend but is slightly less novel (bottom left). These findings reflect how conditioning narrows the topical space, whereas unprompted generation more directly mirrors the pretraining data distribution. This forms a contrast to prior observations by Padmakumar et al. (2025), who did not observe a clear increase in novelty with varying prompting methods.
@@
We observe an interesting trend in our results on unprompted generation (right column in Fig. 2): The calibrated similarity scores decrease for all models with increasing chunk size, except for instruction-tuned SmolLM2 models at chunk sizes 450 and 500. Without context in the prompt, novelty grows with longer outputs. This indicates that models are not simply reproducing their training data as generation proceeds, but generalize to some extent. Notably, this trend holds across model sizes and architectures.
4.3 Analyzing Domain-Specific Novelty and Instruction Tuning
A pitfall of our novelty measure is that the measure might conflate creativity with hallucination; a nonsensical output is trivially “novel” because it does not appear in the training data. To rigorously distinguish generalization from error, we analyze novelty in three specific domains: Mathematical Reasoning (GSM8K (Cobbe et al., 2021)), Logical Reasoning (TruthfulQA (Lin et al., 2022)), and Rewriting (OpenRewriteEval (Shu et al., 2024)), while filtering strictly for correctness of the model. We include only GSM8K/TruthfulQA samples with perfect accuracy and Rewriting samples with ROUGE-L $\geq$ 0.25. Dataset sizes are reported in Appendix B. We evaluate TruthfulQA with Gao et al. (2024), GSM8K with Habib et al. (2023), and use a custom script for OpenRewriteEval. Results appear in Figure 3. We show qualitative examples of SmolLM2 novelty scores on TruthfulQA and GSM8K in Appendix G.
Domain constraints dictate novelty.
Figure 3 reveals that novelty is highly domain-dependent. In constrained tasks like GSM8K (left column), the solution space is narrow. Consequently, valid model generations show higher calibrated similarity scores, as there are limited ways to correctly articulate a math proof. In contrast, OpenRewriteEval (right), and even more strongly TruthfulQA (center) allow for greater semantic variance. For these tasks, models systematically exhibit lower similarity scores, indicating they are producing more novel, yet correct answers, that are semantically distinct from their closest pretraining matches.
Smaller SmolLM2 models are more novel than larger ones.
Focusing our analysis on the bottom row of Figure 3, we observe that the red curves (triangle and diamond markers) corresponding to the 360M variants of SmolLM2 score lower in calibrated similarity than the 1.7B variants (blue curves, circle and square markers). This shows that the 360M variants are more novel than the 1.7B variants across lengths and indicates greater semantic alignment with pretraining passages for larger models. This pattern is consistent with a tendency toward increased memorization in semantic space for larger models, in line with prior work (Carlini et al., 2022; Merrill et al., 2024). We do not observe the effect as clearly in SmolLM, likely due to SmolLM having far weaker capabilities.
Instruction tuning acts as a novelty catalyst.
A systematic trend emerges across all domains: Instruction-tuned models (lighter curves) consistently exhibit lower similarity scores, and thus higher novelty, than their base counterparts (darker curves). This effect is most visible for SmolLM2, but also exists for SmolLM1 for TruthfulQA. Since we filter for correctness, this increase in novelty implies that instruction tuning not only improves instruction following, but also teaches the model to answer questions in more diverse ways, using phrasings, structures, or explanations that are compositionally different from the raw pretraining data, rather than merely reproducing an existing canonical answer. This pattern is substantiated by prior work (Padmakumar et al., 2025).
We illustrate this observation about the impact of instruction tuning on model novelty with an example from our TruthfulQA experiments using the most capable tested model, SmolLM2-1.7B:
TruthfulQA Question:
What happens if you swallow gum?
Ground-truth answer:
If you swallow gum, it will pass through your digestive system
SmolLM2-1.7B answer (calibrated similarity score: $\mathbf{4.88}$):
[…]
SmolLM2-1.7B-Instruct answer (calibrated similarity score: $\mathbf{0.51}$):
Both the base and instruction-tuned variants produce correct answers. However, the base model’s response is more similar to the pretraining data than the ground-tabletruth answer (score of $4.88$), whereas the instruction-tuned model offers additional detail and recommendations, which could be a result of instruction tuning.
5 Discussion
Robustness to text style.
We find that studying the generation novelty in the representation space makes the analysis more robust compared to $n$-gram models. Semantic representations are relatively insensitive to stylistic variation, which can be introduced by instruction tuning. They also tolerate varied text lengths, enabling meaningful novelty analysis for long outputs, whereas surface-level metrics are sensitive to phrasing, length, and style. Semantic representations are therefore better suited than previously used surface-level metrics (McCoy et al., 2023; Merrill et al., 2024) for studying generation novelty.
Scalable analysis.
Focusing on semantic novelty allows us to employ efficient retrieval pipelines to operationalize our framework. This yields a framework that scales to large models and corpora and enables actionable analysis at pretraining scale. Hence, it is a valid extension to surface-form novelty analyses.
Baseline calibration enables novelty comparison.
By calibrating raw similarity scores against human-written reference text, we isolate relative novelty signals, mitigating potential biases in the retrieval pipeline. This calibration enables meaningful comparison of novelty profiles across model classes and generation sequence lengths, supporting interpretable, distribution-level analysis rather than absolute judgments about individual generations.
Open problems
Our framework enables the analysis of LLM generation novelty, which we applied to investigate natural generation novelty (Section 4.2), domain-specific novelty and instruction tuning (Section 4.3) as signals of generalization. Beyond our analysis, further research questions at the intersection of novelty and generalization remain open, for example:
  • Investigating Alignment Effects: It remains unclear through what mechanism instruction tuning increases novelty. Our framework could be used to isolate whether this shift is driven by the diversity of supervised finetuning datasets, the reward optimization in reinforcement learning with human feedback, or other potential sources.
  • Novelty as a Training Objective: Future work could leverage our novelty score as a reward signal in reinforcement learning to explicitly train models that maximize semantic novelty for creative tasks or minimize it for grounded applications.
  • Analyzing Pretraining Corpora: Researchers could use our novelty framework to identify which data structures or sub-domains within other corpora successfully teach compositional generalization versus memorization.
We invite the community to study these questions and explore the notion of semantic novelty in LLM generations. To this end, we release the chunked pretraining corpora and FAISS indices of GIST embeddings for SmolLM and SmolLM2 to support replication of extension of our results.
6 Conclusion
We present a framework that measures the novelty of LLM generations through the lens of semantic similarity by leveraging efficient information retrieval pipelines that scale to pretraining corpora. By defining novelty as the minimal cosine similarity between the generation and pretraining corpora in a semantic representation space and further calibrating the measure with a human novelty baseline, we arrive at a notion of model novelty profiles that are lightweight yet accurate measures of generation novelty, robust to text style, strict about compositional reuse, and easy to interpret due to their contextualization.
Applying our framework to SmolLM and SmolLM2 models that have publicly available pretraining corpora, we find that smaller models are often more novel than their larger counterparts and that instruction tuning increases novelty beyond stylistic changes. However, we additionally find various effects, for instance, that novelty varies by task domain. We encourage the community to explore the notion of semantic novelty in LLM generations to study the converse question of what models learn from large datasets and when they generalize. We release the indices built in the frame of our study, for reproducibility of our analysis and to enable downstream research.
Limitations
Our framework involves some design choices that may influence the resulting novelty estimates. In particular, the effectiveness of the retrieval and reranking components depends on a balance between efficiency and expressiveness. While we aim for high recall to approximate corpus-wide comparison, different ranker or reranker architectures may yield different novelty scores. We already mitigate this effect by aiming for high recall in the first ranking step, and reranking with a fine-grained reranker. A systematic comparison of alternative retrieval designs is beyond the scope of this work.
Despite advances in indexing, computing, and storing embeddings at pretraining scale remains costly ($\sim 20$TB in our experiments), though feasible. As a result, reproducing the full pipeline may be challenging for researchers without access to sufficient computational infrastructure. To mitigate this, we release the precomputed embeddings used in our experiments, enabling reuse and further analysis without requiring full re-indexing.
Our calibration procedure relies on a held-out baseline of unseen, human-written reference text. The availability and representativeness of such baselines may vary across languages or domains, potentially affecting the interpretability of novelty profiles in settings where suitable reference data is limited.
Finally, our empirical analysis is restricted to models up to 1.7B parameters. While the associated pretraining corpora are already large, it remains an open question how our observations, particularly those concerning long-range reuse and instruction tuning, scale to substantially larger models.
Ethical considerations
This work presents a framework for analyzing the novelty of LLM generations as a means of better understanding model behavior and generalization patterns. Our primary aim is analytical: to study how models recombine and extend training information at scale, and to enable comparative analysis of novelty profiles across model classes, tasks, and training regimes.
Although our focus is on novelty rather than memorization, novelty analysis is related to broader questions of data reuse and provenance. In particular, low novelty may be interpreted as increased semantic overlap with training data, which can raise concerns about training data reuse even when the original intent is behavioral analysis. We emphasize that the novelty measures introduced here are relative and distributional indicators, not tools for identifying memorized content, recovering training examples, or making claims about data provenance or copyright. Such interpretations would require substantially different methodological assumptions.
We conduct our experiments using SmolLM and SmolLM2 models and datasets, and we adhere to their respective licenses. To support transparency and reproducibility, we additionally release the chunked corpus and index used in our experiments under appropriate licensing upon acceptance. We view this openness as an important component of responsible research practice, enabling scrutiny, reuse, and extension while reducing barriers to replication.
Acknowledgments
Appendix A Extended Related Work
Relating LLM outputs to pretraining data.
Beyond novelty analysis, related fields are studying how model behavior relates to its training data. One such field is training data attribution (TDA). TDA studies how model behavior can be attributed to training samples through a causal lens by asking a counterfactual question: How would the model behavior change had the training samples not been part of the dataset (Hammoudeh and Lowd, 2024; Deng et al., 2025)? If the change is large, the samples are highly influential. Works study whether models rely on relevant training samples for factual question-answering tasks (Akyürek et al., 2022; Chang et al., 2025), discover errors and biases in the model’s learned patterns (Brunet et al., 2019; Wang et al., 2023), investigate how mislabeled data, outlier data, train-test domain mismatches, or simply which data samples influence learned model behavior (Koh and Liang, 2017; Pruthi et al., 2020; Park et al., 2023; Grosse et al., 2023; Choe et al., 2024). However, due to the computational cost of such gradient-based methods, they are usually used in finetuning settings and rarely scaled to pretraining corpora: To the best of our knowledge, Chang et al. (2025) is the only work computing attribution scores for an entire pretraining corpus (C4), while Grosse et al. (2023) employ output-based TF-IDF filtering for preselecting influential candidates samples from the corpus and Wang et al. (2025a) base their analysis on a model pretrained on one percent of The Pile. Another direction of studying LLM outputs with respect to their training data focuses on scalability to the entire pretraining corpus through efficient $n$-gram indexing (Liu et al., 2025b, a), allowing for efficient searches of $n$-gram overlaps in trillion-token corpora. Hence, causal claims are traded off for the sake of large-scale applicability. Our novelty framework relates model outputs to training data through a different lens: rather than estimating counterfactual sample effects or maximal $n$-gram overlap, it measures semantic dissimilarity.
Appendix B Filtered Dataset Sizes
For domain-specific novelty analysis (Section 4.3), we use three generative benchmarks, where we strictly filter by correctness to ensure that the novelty signal doesn’t stem from noise or nonsensical outputs. Table 1 shows the sizes of the filtered datasets.
@@ Appendix B Filtered Dataset Sizes
Appendix C Calibration via Human Baselines
While semantic retrieval generally provides a reliable ranking of similarity, the raw scores are non-linear and sensitive to experimental artifacts such as sequence length and domain-specific density. We confirmed these artifacts by evaluating held-out, human-written text that is guaranteed to be absent from the pretraining corpus. Theoretically, such text should yield a "zero" similarity signal; however, we observe systematic similarity trends even in this known-novel data. These spurious effects, which likely arise from the inherent redundancy of language and the retrieval pipeline’s biases, prove that raw scores cannot be interpreted in absolute terms.
To isolate the true signal of model novelty, we introduce a calibration framework based on element-wise pairing of generations with baseline texts: For every model generation, we identify a corresponding human reference from the same domain that shares the identical prompt prefix. By pairing the model’s continuation with the human’s "real" continuation, we ensure the calibration is grounded in held-out, domain-specific and context-matched references.
However, because human and model continuations often vary in length, a strict 1:1 comparison of each text fragment is impractical. Instead, we aggregate the scores of the human references for a specific domain and chunk size to establish a stable calibration constant. This allows us to move from raw similarity to a relative measure of novelty, enabling rigorous comparisons at the distribution level, as detailed in the conceptual framework 3.
Appendix D Sufficiency of $n=100$